OpenStudy (anonymous):

Factor completely: 9x^2 + 42x + 49

OpenStudy (anonymous):

Use the (a + b)^2 identity. (3x)^2 + 2(3x)(7) + 7^2 =(3x + 7)^2

OpenStudy (anonymous):

(3x+7)(3x+7)

OpenStudy (anonymous):

(3x + 7)(3x + 7) (3x + 7)(3x - 7) (9x - 7)(9x - 7) (9x + 7)(9x + 7)

OpenStudy (anonymous):

First option.

OpenStudy (anonymous):

(3x-7)^2

OpenStudy (anonymous):

\[\frac{\left( 9x+21 \right)\left( 9x+21 \right)}{9}=\left( 3x+7 \right)^{2}\]

OpenStudy (anonymous):

And Tributized, you made a mistake.

OpenStudy (anonymous):

i know :(